OpenStudy (emily_98):

Ava graphs the function h(x) = x2 + 4. Victor graphs the function g(x) = (x + 4)2. Which statements are true regarding the two graphs? Check all that apply. Ava’s graph is a vertical translation of f(x) = x2. Victor’s graph is a vertical translation of f(x) = x2. Ava’s graph moved 4 units from f(x) = x2 in a positive direction. Victor’s graph moved 4 units from f(x) = x2 in a positive direction. Ava’s graph has a y-intercept of 4. Victor’s graph has a y-intercept of 4.

OpenStudy (anonymous):

lets try Ava’s graph has a y-intercept of 4.

OpenStudy (anonymous):

if \[f(x)=x^2+4\] then \[f(0)=0^2+4=4\] looks good

OpenStudy (anonymous):

some others may be true, but i don't understand what they are asking maybe you do, here are the graphs http://www.wolframalpha.com/input/?i=x^2%2B4,%28x%2B4%29^2+domain+-12..10